# ATLAS Photometry of Interstellar Object 3I/ATLAS

### v1

@@ 2 Observations and Data
StackID MJD filt N m2 m6 m10 m14 dm6 r delta sto $\Phi$
60760-60764o 60762.4563 $o$ 12 24.14 23.74 21.71 21.06 1.88 7.705 7.840 7.29 0.55
60762-60766c 60764.7003 $c$ 12 23.39 22.49 22.78 22.78 0.57 7.628 7.725 7.42 0.55
60791-60792c 60791.5831 $c$ 8 23.27 22.16 21.76 21.94 0.43 6.718 6.340 8.20 0.59
60786-60798o 60792.5380 $o$ 29 23.70 21.61 21.96 21.86 0.19 6.685 6.292 8.17 0.59
60818-60826o 60821.2754 $o$ 17 22.73 20.63 20.14 19.90 0.07 5.714 4.897 6.43 0.51
60817-60828c 60823.0204 $c$ 17 22.83 21.29 20.86 21.25 0.17 5.655 4.818 6.22 0.50
04a60843 60843.2952 $o$ 4 20.98 19.43 18.92 18.81 0.08 4.973 3.976 2.55 0.27
05r60845 60845.0836 $w$ 4 20.86 19.31 18.81 18.66 0.05 4.913 3.910 2.14 0.23
05r60848 60848.0824 $w$ 4 20.88 19.13 18.86 18.76 0.04 4.812 3.802 1.47 0.17
04a60850 60850.4007 $o$ 4 20.47 18.62 18.23 18.09 0.04 4.735 3.722 1.09 0.13
05r60851 60851.0416 $w$ 3 21.12 19.18 18.70 18.52 0.04 4.714 3.700 1.03 0.13
03a60851 60851.0812 $c$ 5 21.72 19.65 19.06 18.73 0.06 4.712 3.698 1.03 0.13
04a60853 60853.2509 $o$ 4 20.26 18.31 17.86 17.80 0.05 4.640 3.626 1.10 0.13
02a60854 60854.4585 $c$ 4 21.11 19.12 18.61 18.47 0.05 4.599 3.587 1.30 0.16
05r60855 60855.0381 $w$ 4 20.45 18.57 18.17 17.98 0.03 4.580 3.569 1.42 0.17
03a60856 60855.9503 $c$ 4 21.02 19.03 18.52 18.31 0.02 4.550 3.540 1.64 0.19
04a60857 60857.2336 $o$ 4 20.10 18.34 17.97 17.84 0.02 4.507 3.500 1.99 0.22
05r60858 60858.0347 $w$ 4 20.28 18.50 18.23 18.16 0.02 4.480 3.475 2.23 0.24
04a60858 60858.2348 $o$ 4 20.45 18.45 17.95 17.78 0.04 4.473 3.469 2.30 0.25
03a60859 60858.9859 $c$ 4 20.72 18.81 18.35 18.20 0.02 4.449 3.447 2.53 0.27
04a2300 60860.9796 $o$ 18 20.24 18.32 17.92 17.73 0.02 4.382 3.388 3.18 0.32
05r60861 60861.0135 $w$ 4 20.31 18.47 18.06 17.98 0.03 4.381 3.387 3.19 0.32
04a0000 60861.0213 $o$ 21 20.15 18.27 17.83 17.69 0.02 4.381 3.387 3.19 0.32
04a0100 60861.0620 $o$ 18 20.14 18.29 17.85 17.72 0.02 4.379 3.386 3.21 0.32
04a0200 60861.1044 $o$ 16 20.21 18.27 17.87 17.75 0.02 4.378 3.384 3.22 0.32
04a0300 60861.1450 $o$ 17 20.25 18.30 17.87 17.72 0.02 4.377 3.383 3.23 0.32
04a0400 60861.1875 $o$ 18 20.20 18.24 17.82 17.65 0.02 4.375 3.382 3.25 0.32
04a0500 60861.2296 $o$ 16 20.21 18.28 17.85 17.72 0.02 4.374 3.381 3.26 0.32
04a0600 60861.2717 $o$ 14 20.26 18.34 17.87 17.69 0.02 4.372 3.380 3.28 0.32
02a0600 60861.2733 $o$ 19 20.12 18.29 17.87 17.68 0.02 4.372 3.380 3.28 0.32
02a0700 60861.3128 $o$ 23 20.27 18.31 17.91 17.76 0.01 4.371 3.378 3.29 0.33
04a0700 60861.3128 $o$ 14 20.25 18.37 17.93 17.80 0.02 4.371 3.378 3.29 0.33
04a0800 60861.3508 $o$ 12 20.22 18.38 17.88 17.64 0.03 4.370 3.377 3.30 0.33
02a0800 60861.3556 $o$ 25 20.27 18.34 17.91 17.76 0.01 4.369 3.377 3.31 0.33
02a0900 60861.3953 $o$ 24 20.30 18.36 17.94 17.79 0.01 4.368 3.376 3.32 0.33
02a1000 60861.4382 $o$ 22 20.31 18.34 17.89 17.73 0.02 4.367 3.375 3.33 0.33
02a1100 60861.4789 $o$ 20 20.22 18.39 17.93 17.72 0.02 4.365 3.374 3.35 0.33
02a1200 60861.5204 $o$ 17 20.41 18.40 17.91 17.73 0.02 4.364 3.373 3.36 0.33
04a60862 60862.1202 $o$ 4 20.37 18.42 17.84 17.62 0.03 4.344 3.355 3.57 0.35
05r60868 60868.0385 $w$ 4 19.77 17.99 17.68 17.57 0.04 4.148 3.199 5.72 0.47
04a60868* 60868.1667 $o$ 4 20.00 18.14 17.64 17.54 0.04 4.143 3.195 5.78 0.48
03a60869 60868.9227 $o$ 4 20.10 18.08 17.57 17.38 0.05 4.119 3.177 6.06 0.49
04a60869 60869.2095 $o$ 4 19.95 18.08 17.60 17.43 0.03 4.109 3.170 6.17 0.50
04a60870 60870.1940 $o$ 7 19.91 18.05 17.62 17.39 0.03 4.076 3.146 6.56 0.51
05r60871 60870.9906 $w$ 4 20.23 18.32 17.90 17.72 0.02 4.050 3.127 6.87 0.53
03a60871 60871.0092 $o$ 7 20.08 18.19 17.78 17.63 0.03 4.049 3.127 6.88 0.53
03a60872 60872.0120 $o$ 7 19.90 18.07 17.62 17.39 0.02 4.016 3.104 7.27 0.55
03a60873 60872.8108 $o$ 2 19.95 18.07 17.64 17.46 0.02 3.990 3.086 7.58 0.56
05r60874 60873.9749 $w$ 4 20.28 18.34 17.85 17.63 0.02 3.951 3.060 8.06 0.58
04a60874 60874.1477 $c$ 4 20.17 18.26 17.84 17.67 0.02 3.945 3.056 8.13 0.58
03a60875 60874.8838 $o$ 4 19.92 17.94 17.49 17.29 0.01 3.922 3.041 8.42 0.60
03a60876 60875.8850 $o$ 5 19.98 17.89 17.35 17.15 0.02 3.888 3.020 8.83 0.61
03a60877 60876.7542 $o$ 2 19.77 17.80 17.26 17.06 0.02 3.860 3.002 9.18 0.63
05r60877 60876.9726 $w$ 4 19.77 17.92 17.55 17.37 0.01 3.852 2.998 9.28 0.63
04a60877 60877.1447 $c$ 4 19.98 18.11 17.65 17.44 0.03 3.847 2.995 9.34 0.63
04a60878 60878.1331 $c$ 4 20.06 18.09 17.64 17.48 0.02 3.814 2.975 9.75 0.65
03a60880 60879.9488 $o$ 4 19.58 17.66 17.19 17.02 0.01 3.755 2.941 10.50 0.68
05r60880 60879.9681 $w$ 4 20.30 18.22 17.68 17.44 0.01 3.754 2.941 10.51 0.68
04a60881 60881.1325 $c$ 4 19.82 17.93 17.50 17.32 0.01 3.716 2.920 10.99 0.70
03a60883* 60882.8439 $o$ 4 19.46 17.53 17.06 16.88 0.01 3.660 2.890 11.70 0.72
05r60883 60882.9651 $w$ 8 20.06 18.06 17.52 17.28 0.01 3.656 2.888 11.75 0.72
03a60884 60883.8467 $o$ 4 19.43 17.47 17.01 16.83 0.01 3.627 2.874 12.11 0.74
04a60885 60885.0847 $c$ 4 19.71 17.84 17.42 17.21 0.01 3.587 2.854 12.62 0.75
05r60886 60885.9606 $w$ 4 19.44 17.63 17.24 17.10 0.01 3.558 2.841 12.98 0.77
02a60886 60886.3258 $o$ 4 19.29 17.33 16.85 16.67 0.01 3.546 2.835 13.13 0.77
05r60889 60888.9049 $w$ 4 19.66 17.72 17.19 16.98 0.01 3.462 2.799 14.17 0.81
04a60894 60894.1118 $o$ 4 19.00 17.08 16.58 16.33 0.03 3.294 2.735 16.21 0.88
03a60895 60894.8868 $o$ 4 19.02 17.07 16.49 16.35 0.03 3.269 2.727 16.50 0.89
05r60896 60895.9044 $w$ 4 19.19 17.29 16.76 16.61 0.03 3.236 2.716 16.88 0.90
05r60900 60899.8984 $w$ 4 19.14 17.20 16.74 16.57 0.01 3.109 2.679 18.30 0.95
03a60905 60904.8024 $o$ 4 18.84 16.80 16.26 16.03 0.01 2.953 2.642 19.86 1.00
05r60908 60907.8897 $w$ 4 18.96 16.97 16.47 16.28 0.01 2.856 2.623 20.73 1.02
04a60914 60914.0567 $o$ 4 18.44 16.40 15.88 15.65 0.01 2.666 2.593 22.11 1.07
02a60916 60916.2574 $o$ 4 18.36 16.34 15.83 15.59 0.01 2.599 2.585 22.48 1.08
03a60917 60916.7302 $o$ 4 18.41 16.37 15.82 15.59 0.01 2.584 2.583 22.55 1.08
The magnitudes from Teide ($w$ band) are subject to revision, but we list the $w$ band results in Table LABEL:tab:phot to provide knowledge of when ATLAS $w$ band exposures exist. Teide uses a complex system of 16 Celestron RASA-11 telescopes co-aligned in groups of 4, and it takes 10 sec sub-exposures. The 12 sub-exposures for each field are all aligned and stacked into a single image, but the sub-exposures do not overlap perfectly, and the PSF can be quite complex because of wind buffeting and focus. In particular, 3I/ATLAS was closer than 4′ to the edge of the images on 05r60868 and 05r60883 and lack of coverage by all of the 16 sub-exposures may cause the photometric calibration to be in error. The reduction pipeline for Teide is also still evolving, so if it proves to be important it is possible to re-reduce the Teide images of 3I/ATLAS to improve the accuracy of the $w$ band photometry, perhaps by using a subset of sub-exposures.
A way to judge the significance of the early detections is to ask how frequently a $\times6$″ aperture in the stacked detection is brighter than the one that is selected by the known orbit of 3I/ATLAS. For the first stack epoch, encompassing MJD 60760–60766, the orbit selected pixel is exceeded 1.2% of the time for both filters, 2.2% of the time for the $c$ filter, and 4.4% of the time for the $o$ filter. Gaussian statistics would call these $2\sigma$ detections”. For the second epoch, MJD 60786–60798, the orbit selected pixel is exceeded 0.04% of the time for both filters combined, 1.3% of the time for the $c$ filter, and 0.1% of the time for the $o$ filter. Gaussian statistics would call these $3\sigma$ detections”. In fact the significance is better than this because we could condition consideration of false alarms by whether there is a bright blob in the stack before differencing and disregard those pixels (none of the early detection stacks has a bright blob at the location from the orbit). The uncertainties in Table LABEL:tab:phot are therefore fairly reliable indicators of false alarm probability as well as measurement uncertainty. After the first two epochs the orbit centered pixels are much brighter than any other among $10^{4}$ trials and plainly visible to the eye.
Figure 3 illustrates the evolution of the brightness of 3I/ATLAS in a pair of graphs. Left are the magnitudes as a function of time, right are the magnitudes corrected by distance from Sun $r$, Earth $\delta$, and angular $G=0.15$ phase function to absolute $H$ magnitude: the brightness of a fully illuminated object if at 1 au distance from both Sun and Earth (not physically realizable).
@@
Figure 3: Magnitudes of 3I/ATLAS in a $\times6$″ aperture are shown on the left as a function of time. Red is $o$ band, blue is $c$ band, magenta is $w$ band from Teide. On the right the magnitudes are corrected for geometrical illumination and distance, $5\log(r\delta)$, and a standard $G=0.15$ phase function, i.e. an absolute $H$ magnitude in the observed bandpass. The actual dependence of brightness on phase angle is unknown, but most likely does not deviate by a $G=0.15$ function for asteroids by more than $\sim 0.1$ mag over this range of sto.
[…]

### v3

@@ 2 Observations and Data
StackID MJD filt N m2 m6 m10 m14 dm6 r Delta sto HM
60760-60764o 60762.4563 $o$ 12 24.14 23.74 21.71 21.06 1.88 7.705 7.840 7.29 0.30
60762-60766c 60764.7003 $c$ 12 23.39 22.49 22.78 22.78 0.57 7.628 7.725 7.42 0.31
60791-60792c 60791.5831 $c$ 8 23.27 22.16 21.76 21.94 0.43 6.718 6.340 8.20 0.34
60786-60798o 60792.5380 $o$ 29 23.70 21.61 21.96 21.86 0.19 6.685 6.292 8.17 0.34
60817-60828o 60821.2754 $o$ 17 22.73 20.63 20.14 19.90 0.07 5.714 4.897 6.43 0.27
60818-60826c 60823.0204 $c$ 17 22.83 21.29 20.86 21.25 0.17 5.655 4.818 6.22 0.26
04a60843 60843.2952 $o$ 4 20.98 19.43 18.92 18.81 0.08 4.973 3.976 2.55 0.11
05r60845 60845.0836 $w$ 4 20.86 19.31 18.81 18.66 0.05 4.913 3.910 2.14 0.09
05r60848 60848.0824 $w$ 4 20.88 19.13 18.86 18.76 0.04 4.812 3.802 1.47 0.07
04a60850 60850.4007 $o$ 4 20.47 18.62 18.23 18.09 0.04 4.735 3.722 1.09 0.05
05r60851 60851.0416 $w$ 3 21.12 19.18 18.70 18.52 0.04 4.714 3.700 1.03 0.05
03a60851 60851.0812 $c$ 5 21.72 19.65 19.06 18.73 0.06 4.712 3.698 1.03 0.05
04a60853 60853.2509 $o$ 4 20.26 18.31 17.86 17.80 0.05 4.640 3.626 1.10 0.05
02a60854 60854.4585 $c$ 4 21.11 19.12 18.61 18.47 0.05 4.599 3.587 1.30 0.06
05r60855 60855.0381 $w$ 4 20.45 18.57 18.17 17.98 0.03 4.580 3.569 1.42 0.06
03a60856 60855.9503 $c$ 4 21.02 19.03 18.52 18.31 0.02 4.550 3.540 1.64 0.07
04a60857 60857.2336 $o$ 4 20.10 18.34 17.97 17.84 0.02 4.507 3.500 1.99 0.09
05r60858 60858.0347 $w$ 4 20.28 18.50 18.23 18.16 0.02 4.480 3.475 2.23 0.10
04a60858 60858.2348 $o$ 4 20.45 18.45 17.95 17.78 0.04 4.473 3.469 2.30 0.10
03a60859 60858.9859 $c$ 4 20.72 18.81 18.35 18.20 0.02 4.449 3.447 2.53 0.11
04a2300 60860.9796 $o$ 18 20.24 18.32 17.92 17.73 0.02 4.382 3.388 3.18 0.14
05r60861 60861.0135 $w$ 4 20.31 18.47 18.06 17.98 0.03 4.381 3.387 3.19 0.14
04a0000 60861.0213 $o$ 21 20.15 18.27 17.83 17.69 0.02 4.381 3.387 3.19 0.14
04a0100 60861.0620 $o$ 18 20.14 18.29 17.85 17.72 0.02 4.379 3.386 3.21 0.14
04a0200 60861.1044 $o$ 16 20.21 18.27 17.87 17.75 0.02 4.378 3.384 3.22 0.14
04a0300 60861.1450 $o$ 17 20.25 18.30 17.87 17.72 0.02 4.377 3.383 3.23 0.14
04a0400 60861.1875 $o$ 18 20.20 18.24 17.82 17.65 0.02 4.375 3.382 3.25 0.14
04a0500 60861.2296 $o$ 16 20.21 18.28 17.85 17.72 0.02 4.374 3.381 3.26 0.14
04a0600 60861.2717 $o$ 14 20.26 18.34 17.87 17.69 0.02 4.372 3.380 3.28 0.14
02a0600 60861.2733 $o$ 19 20.12 18.29 17.87 17.68 0.02 4.372 3.380 3.28 0.14
02a0700 60861.3128 $o$ 23 20.27 18.31 17.91 17.76 0.01 4.371 3.378 3.29 0.14
04a0700 60861.3128 $o$ 14 20.25 18.37 17.93 17.80 0.02 4.371 3.378 3.29 0.14
04a0800 60861.3508 $o$ 12 20.22 18.38 17.88 17.64 0.03 4.370 3.377 3.30 0.14
02a0800 60861.3556 $o$ 25 20.27 18.34 17.91 17.76 0.01 4.369 3.377 3.31 0.14
02a0900 60861.3953 $o$ 24 20.30 18.36 17.94 17.79 0.01 4.368 3.376 3.32 0.14
02a1000 60861.4382 $o$ 22 20.31 18.34 17.89 17.73 0.02 4.367 3.375 3.33 0.14
02a1100 60861.4789 $o$ 20 20.22 18.39 17.93 17.72 0.02 4.365 3.374 3.35 0.15
02a1200 60861.5204 $o$ 17 20.41 18.40 17.91 17.73 0.02 4.364 3.373 3.36 0.15
04a60862 60862.1202 $o$ 4 20.37 18.42 17.84 17.62 0.03 4.344 3.355 3.57 0.16
05r60868* 60868.0385 $w$ 4 19.77 17.99 17.68 17.57 0.04 4.148 3.199 5.72 0.24
04a60868 60868.1667 $o$ 4 20.00 18.14 17.64 17.54 0.04 4.143 3.195 5.78 0.25
03a60869 60868.9227 $o$ 4 20.10 18.08 17.57 17.38 0.05 4.119 3.177 6.06 0.26
04a60869 60869.2095 $o$ 4 19.95 18.08 17.60 17.43 0.03 4.109 3.170 6.17 0.26
04a60870 60870.1940 $o$ 7 19.91 18.05 17.62 17.39 0.03 4.076 3.146 6.56 0.28
05r60871 60870.9906 $w$ 4 20.23 18.32 17.90 17.72 0.02 4.050 3.127 6.87 0.29
03a60871 60871.0092 $o$ 7 20.08 18.19 17.78 17.63 0.03 4.049 3.127 6.88 0.29
03a60872 60872.0120 $o$ 7 19.90 18.07 17.62 17.39 0.02 4.016 3.104 7.27 0.30
03a60873 60872.8108 $o$ 2 19.95 18.07 17.64 17.46 0.02 3.990 3.086 7.58 0.32
05r60874 60873.9749 $w$ 4 20.28 18.34 17.85 17.63 0.02 3.951 3.060 8.06 0.33
04a60874 60874.1477 $c$ 4 20.17 18.26 17.84 17.67 0.02 3.945 3.056 8.13 0.34
03a60875 60874.8838 $o$ 4 19.92 17.94 17.49 17.29 0.01 3.922 3.041 8.42 0.35
03a60876 60875.8850 $o$ 5 19.98 17.89 17.35 17.15 0.02 3.888 3.020 8.83 0.36
03a60877 60876.7542 $o$ 2 19.77 17.80 17.26 17.06 0.02 3.860 3.002 9.18 0.38
05r60877 60876.9726 $w$ 4 19.77 17.92 17.55 17.37 0.01 3.852 2.998 9.28 0.38
04a60877 60877.1447 $c$ 4 19.98 18.11 17.65 17.44 0.03 3.847 2.995 9.34 0.38
04a60878 60878.1331 $c$ 4 20.06 18.09 17.64 17.48 0.02 3.814 2.975 9.75 0.40
03a60880 60879.9488 $o$ 4 19.58 17.66 17.19 17.02 0.01 3.755 2.941 10.50 0.42
05r60880 60879.9681 $w$ 4 20.30 18.22 17.68 17.44 0.01 3.754 2.941 10.51 0.42
04a60881 60881.1325 $c$ 4 19.82 17.93 17.50 17.32 0.01 3.716 2.920 10.99 0.44
03a60883 60882.8439 $o$ 4 19.46 17.53 17.06 16.88 0.01 3.660 2.890 11.70 0.47
05r60883* 60882.9651 $w$ 8 20.06 18.06 17.52 17.28 0.01 3.656 2.888 11.75 0.47
03a60884 60883.8467 $o$ 4 19.43 17.47 17.01 16.83 0.01 3.627 2.874 12.11 0.48
04a60885 60885.0847 $c$ 4 19.71 17.84 17.42 17.21 0.01 3.587 2.854 12.62 0.50
05r60886 60885.9606 $w$ 4 19.44 17.63 17.24 17.10 0.01 3.558 2.841 12.98 0.51
02a60886 60886.3258 $o$ 4 19.29 17.33 16.85 16.67 0.01 3.546 2.835 13.13 0.51
05r60889 60888.9049 $w$ 4 19.66 17.72 17.19 16.98 0.01 3.462 2.799 14.17 0.55
04a60894 60894.1118 $o$ 4 19.00 17.08 16.58 16.33 0.03 3.294 2.735 16.21 0.61
03a60895 60894.8868 $o$ 4 19.02 17.07 16.49 16.35 0.03 3.269 2.727 16.50 0.62
05r60896 60895.9044 $w$ 4 19.19 17.29 16.76 16.61 0.03 3.236 2.716 16.88 0.63
05r60900 60899.8984 $w$ 4 19.14 17.20 16.74 16.57 0.01 3.109 2.679 18.30 0.68
03a60905 60904.8024 $o$ 4 18.84 16.80 16.26 16.03 0.01 2.953 2.642 19.86 0.72
05r60908 60907.8897 $w$ 4 18.96 16.97 16.47 16.28 0.01 2.856 2.623 20.73 0.74
04a60914 60914.0567 $o$ 4 18.44 16.40 15.88 15.65 0.01 2.666 2.593 22.11 0.78
02a60916 60916.2574 $o$ 4 18.36 16.34 15.83 15.59 0.01 2.599 2.585 22.48 0.79
03a60917 60916.7302 $o$ 4 18.41 16.37 15.82 15.59 0.01 2.584 2.583 22.55 0.79
The magnitudes from Teide ($w$ band) are subject to revision, but we list the $w$ band results in Table LABEL:tab:phot to provide knowledge of when ATLAS $w$ band exposures exist. Teide uses a complex system of 16 Celestron RASA-11 telescopes co-aligned in groups of 4, and it takes 10 sec sub-exposures. The 12 sub-exposures for each field are all aligned and stacked into a single image, but the telescopes do not overlap perfectly, and the PSF can be quite complex because of wind buffeting and focus. In particular, 3I/ATLAS was closer than 4′ to the edge of the images on 05r60868 and 05r60883 and lack of coverage by all of the 16 sub-exposures may cause the photometric calibration to be in error. The reduction pipeline for Teide is also still evolving, so if it proves to be important it is possible to re-reduce the Teide images of 3I/ATLAS to improve the accuracy of the $w$ band photometry, perhaps by using a subset of sub-exposures.
A way to judge the significance of the early detections is to ask how frequently a $\times6$″ aperture in the stacked image is brighter than the one that is selected by the known orbit of 3I/ATLAS. For the first stack epoch, encompassing MJD 60760–60766, the orbit selected pixel is exceeded 1.2% of the time for both filters, 2.2% of the time for the $c$ filter, and 4.4% of the time for the $o$ filter. Gaussian statistics would call these $2\sigma$ detections”. For the second epoch, MJD 60786–60798, the orbit selected pixel is exceeded 0.04% of the time for both filters combined, 1.3% of the time for the $c$ filter, and 0.1% of the time for the $o$ filter. Gaussian statistics would call these $3\sigma$ detections”. In fact the significance is better than this because we could condition consideration of false alarms by whether there is a bright blob in the stack before differencing and disregard those pixels (none of the early detection stacks has a bright blob at the location from the orbit). The uncertainties in Table LABEL:tab:phot are therefore fairly reliable indicators of false alarm probability as well as measurement uncertainty. After the first two epochs the orbit centered pixels are much brighter than any other among $10^{4}$ trials and plainly visible to the eye.
Figure 3 illustrates the evolution of the brightness of 3I/ATLAS in a pair of graphs. Left are the magnitudes as a function of time, right are the magnitudes corrected by distance from Sun $r$, Earth $\Delta$, and angular phase function to absolute $H(1,1,0)$ magnitude: the brightness of a fully illuminated object if at 1 au distance from both Sun and Earth (not physically realizable).
@@
Figure 3: Magnitudes of 3I/ATLAS in a $\times6$″ aperture are shown on the left as a function of time. Red is $o$ band, blue is $c$ band, magenta is $w$ band from Teide. On the right the magnitudes are corrected for geometrical illumination and distance, $5\log(r\Delta)$, and the Schleicher HM phase function, i.e. an absolute $H$ magnitude in the observed bandpass that shows the physical evolution of 3I/ATLAS with time.
Of course it should be understood that most of the light at all times was coming from the extended coma, not the nucleus of this comet, evident in the brightening in successive apertures. Therefore application of the HM phase function for these evolving coma particles to construct the $H$ magnitude may be misleading. As seen in the table, the HM phase function is very nearly 0.04 mag/deg for the observed $4^{\circ}<\hbox{sto}<23^{\circ}$, but this can be revised as desired for a given physical model. Similarly the volume encompassed by the apertures is a 2 dimensional pencil beam through at 3 dimensional coma and the geometrical correction must be interpreted with this in mind.